How do you get 0.501 M with 41.0 mL and 3.41M and 279mL


1
Expert's answer
2021-05-06T07:39:09-0400

Using formula of resultant molality-


"m=\dfrac{m_1v_1+m_2v_2}{v_1+v_2}"


"=\dfrac{0.501\times 41+3.41\times 279}{41+279}\\[9pt]=\dfrac{971.931}{320}=3.07M"